OpenStudy (anonymous):

lim x to 1 of (sin (x*pi))/(x-1)

OpenStudy (anonymous):

Ahhhh without the use of L'opithals rule!!

OpenStudy (mertsj):

Have to ask the smart people.

OpenStudy (anonymous):

make the substitution \(u=x-1\) the limits change as \(x\to1\), \(x-1\to0\implies u\to0\)

OpenStudy (anonymous):

then it becomes a standard limit!

OpenStudy (anonymous):

It will become lim u to 0 of sen((u+1)pi)/u right?

OpenStudy (anonymous):

yes. but then you can expand that \[\sin(u+1)\pi=\sin(u\pi+\pi)=-\sin(u\pi)\]

OpenStudy (anonymous):

I'm really sorry. After that u'll get - sin(u*pi)/u then what? :(

OpenStudy (anonymous):

that is a standard form remember?

OpenStudy (anonymous):

Lim u to 0 of sin u/ u is 1 but in here we have pi multipling inside the factor not outside (sin (u*pi) not pi*sin(u) )right?

OpenStudy (anonymous):

exactly.. so, multiply and divide by "pi" this will create a "pi*x" at the denominator

OpenStudy (anonymous):

Omg, now i see it. Sorry man thx a lot!!!

OpenStudy (anonymous):

no problemo.